Circular No. 7340

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

QY PERSEI J. A. Mattei, AAVSO, informs us that this suspected SU UMa-type

variable is undergoing what may be a superoutburst, as indicated by the following visual magnitude estimates: Dec. 19.972 UT, [14.9 (G. Poyner, W. Midlands, England); 28.146, 13.5 (M. Simonsen, Macomb, MI); 28.166, 13.6 (Simonsen); 29.731, 13.9 (E. Muyllaert, Oostende, Belgium); 29.828, 13.8 (C. Jones, Essex, England); 29.901, 13.8 (H. McGee, Surrey, England). Mattei notes that high-speed photometry to look for superhumps is recommended. The last confirmed outburst for QY Per in the AAVSO archive occurred on 1995 Aug. 21, at peak mag 14.1. Additional unfiltered CCD magnitudes: Dec. 23.190, [17.5 (P. Schmeer, Bischmisheim, Germany); 28.168, 13.9 (Schmeer); 29.887, 14.2 (G. Masi, Ceccano, Italy). Masi also provides the following precise position for QY Per: α = 3^{h}15^{m}36^{s}.85, δ = +42^{o}28′15′′.3 (equinox 2000.0).

SUPERNOVA 1999ga IN NGC 2442 M. Salvo, L. Rizzi, A. Pastorello, and M. Turatto, Padova Astronom-

ical Observatory, report: “A spectrum (370–910 nm; resolution 1.1 nm) of SN 1999ga (cf. IAUC 7316) was taken on Dec. 29.26 UT with the 1.54-m Danish telescope (+ DFOSC) at the European Southern Observatory. Preliminary reduction of the spectrum shows lines of H, Ca II, [Ca II], Na I D, and Fe II, indicating that the supernova is of type II. We therefore confirm the suggestion by Rubenstein (IAUC 7319), which was based on the supernova’s location inside the parent galaxy, although our spectrum shows no strong evidence of interstellar absorption. The supernova seems rather evolved, and the line emissions dominate over the absorptions. Indeed, an automated comparison with all the spectra in our supernovae archive shows the best fit to be with the spectrum of SN 1992H at 143 days after maximum (Clocchiatti et al. 1996, A.J. 111, 1286). The shape of the continuum is also similar, indicating that the two supernovae suffered comparable reddening. The strongest emission line (Hα) is flat-topped, suggesting that the supernova ejecta is interacting with the circumstellar medium.”

SUPERNOVA 1999gp IN UGC 1993 K. Krisciunas and K. Loomis report preliminary infrared photometry

for SN 1999gp, obtained as noted for SN 1999go on IAUC 7339: Dec. 27.07 UT, J = 15.92 ± 0.07, K′ = 15.05 ± 0.08. They add that SN 1999gp is not the brighter foreground star located just north of the galaxy’s core.

1999 December 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green